OpenStudy (anonymous):

A rail is 10 m long. It must be supported at least once every 3 m, beginning with a support at the far left end. Counting the first support, at how many points must a support be added to the rail? 2 points 3 points 4 points 5 points

OpenStudy (anonymous):

It's 2 points. What do you think?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

If you took ten and separated it into groups of three, one would be left over. That's how I got my answer. :)

OpenStudy (anonymous):

sounds good

OpenStudy (anonymous):

Serenity, Lucy, and Piper hiked for 4 hours. They hiked 1\frac{3}{4}miles each hour for the first three hours. In the fourth hour, they hiked \frac{3}{10}of the total distance that they hiked during the first three hours. What was the total distance the girls \frac{31}{40} \frac{5}{6} \frac{5}{6} \frac{33}{40}